2m-nx=x+4 solve for x work needed
Found 2 solutions by solver91311, checkley77:
Answer by solver91311(24713)


Add -x to both sides

Add -2m to both sides

Multiply both sides by -1

Factor x out of the LHS

Divide both sides by n + 1


John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by checkley77(12844)

2m-nx=x+4 add nx
x+nx=2m-4 factor x
x(n-1)=2m-4 divide by (n-1)
x=(2m-4)/(n-1) ans.